OpenStudy (anonymous):

Fan or Medal The equation, y = 12841(1.01)x , represents the average cost of an education at a private college, after 1990. What is the average cost in 1990 (x = 0)? $12,969 $12,841 $13,122 $13,099

OpenStudy (blurbendy):

12841, because 1 year after 1990, the avg tuition would be 12,969

jimthompson5910 (jim_thompson5910):

you plug in x = 0 and evaluate to get your answer

OpenStudy (anonymous):

12969.41??

OpenStudy (blurbendy):

12969.41 is the average tuition after 1 year

OpenStudy (blurbendy):

in 1990 it's simply going to be 12841

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

y is the average cost each year

jimthompson5910 (jim_thompson5910):

so you just plug in x = 0 and you evaluate

OpenStudy (blurbendy):

but then dont you get 0?

jimthompson5910 (jim_thompson5910):

the equation is y = 12841(1.01)^x the symbol ^ didn't make it when it was pasted

OpenStudy (blurbendy):

oh! lol ok

jimthompson5910 (jim_thompson5910):

cost is usually modeled by exponential functions

OpenStudy (anonymous):

can one of you draw a example?

OpenStudy (anonymous):

@jim_thompson5910 @blurbendy

jimthompson5910 (jim_thompson5910):

exponential functions look like this 

jimthompson5910 (jim_thompson5910):

ie the cost grows at a slow rate at first, but that growth rate increases over time

OpenStudy (anonymous):

I understand that but how do you work the problem and why do you put your number where you put them @jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = 12841(1.01)^x y = 12841(1.01)^0 ... plug in x = 0 y = 12841(1) y = 12841

jimthompson5910 (jim_thompson5910):

so that's all there is to it really

OpenStudy (anonymous):

okay this is not part of the question but what if x=50?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

plug in x = 50 and evaluate

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

how would I mutilpy out x, do I do it like a expont or do I just times it or what do I do is my question really? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you use a calculator, so y = 12841(1.01)^x y = 12841(1.01)^(50) ... plug in x = 50 y = 12841(1.64463182184389) ... use a calculator y = 21,118.7172242973 ... use a calculator y = 21,118.72 ... round to the nearest penny

jimthompson5910 (jim_thompson5910):

so the average cost goes from $12,841 in 1990 to $21,118.72 in 2040 (50 years later)

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yea okay so what if they give me the Y but not the x? how do I work this one?

jimthompson5910 (jim_thompson5910):

you plug in the given value for y, then solve for x

jimthompson5910 (jim_thompson5910):

you would use logs to do so

OpenStudy (anonymous):

how would I slove for x, if you don't mind please show me the way you did step by step I understood that better

jimthompson5910 (jim_thompson5910):

what value of y are they giving you?

OpenStudy (anonymous):

they didn't give me one but we can use 11 or just choose a number

jimthompson5910 (jim_thompson5910):

so why do you want to solve for something they didn't give you?

OpenStudy (anonymous):

so I know how to work it when it comes time?

jimthompson5910 (jim_thompson5910):

are they asking you to solve for x? without giving the value of y

jimthompson5910 (jim_thompson5910):

i gotcha

jimthompson5910 (jim_thompson5910):

ok pick any cost value between 10 thousand and 20 thousand

OpenStudy (anonymous):

11,356 is that okay?

jimthompson5910 (jim_thompson5910):

sure